Can you help me solve this? If 15 is added to a number, the result is 56 less than twice the number. Find the number.
Answer by Theo(13342)

x = the number
x + 15 = 2x - 56
subtract x from both sides of the equation and add 56 to both sides of the equation to get:
x = 71
that should be your answer.
add 15 to that number and you get 71 + 15 = 86.
take 2 times the number and subtract 56 from it to get 2 * 71 - 56 = 142 - 56 = 86
number checks out so the answer is good.